How to find the equation of a line which passes through the point of intersection of the of the lines x - 3y =4 and 3x + y =0 and are relatively parallel and perpendicular to the line 3x + 4y =0


1
Expert's answer
2021-03-01T17:04:11-0500

Find the point of intersection of the of the lines x - 3y =4 and 3x + y =0 


"y=\dfrac{1}{3}x-\dfrac{4}{3}\ and\ y=-3x"

"\dfrac{1}{3}x-\dfrac{4}{3}=-3x"

"x=\dfrac{2}{5}"

"y=-\dfrac{6}{5}"

"Point\ \big(\dfrac{2}{5}, -\dfrac{6}{5}\big)"

Parallel to the line 3x + 4y =0


"slope=m_1=-\dfrac{3}{4}"

"y-(-\dfrac{6}{5})=-\dfrac{3}{4}(x-\dfrac{2}{5})"

"3x+4y=-\dfrac{18}{5}"

Perpendicular to the line 3x + 4y =0


"slope=m_2=\dfrac{4}{3}"

"y-(-\dfrac{6}{5})=\dfrac{4}{3}(x-\dfrac{2}{5})"

"4x-3y=\dfrac{26}{5}"